# Ergodicity for locally monotone stochastic evolution equations with Lévy noise

### v1

@@ 3.2. Hypotheses on the noise
The noise coefficient $B:H\to H$ is linear, symmetric, bounded, and positive definite, and satisfies
(3.3) $B\in L_{2}(H),$
where $\|B\|_{2}:=\|B\|_{L_{2}(H)}$. The intensity measure $\pi$ of $N$ satisfies (2.2). The strongly measurable jump coefficient $G:Z\to H$ satisfies
$\mathfrak{g}_{0}:=\int_{D^{c}}\|G(z)\|_{H}^{2}\,\pi(\mathrm{d}z)<\infty,$
@@
$\mathfrak{g}_{\beta}:=\int_{D^{c}}\|G(z)\|_{H}^{\beta+2}\,\pi(\mathrm{d}z)<\infty,$
where $\beta\geq 0$ is as in Hypothesis (C).
Hypothesis (G) ($V$-Regularity of the noise).
For any $T>0$, the driving process of the noise
(3.4) $L_{t}:=B\,W_{t}+\int_{D^{c}}G(z)\,\widetilde{N}(\{t\},\mathrm{d}z),\quad t\in[0,T],$
where $(W_{t})_{t\in[0,T]}$ and $\left(\widetilde{N}(\{t\}\times U)\right)_{t\in[0,T],\;U\in\mathcal{Z}}$ are independent, has a representation
(3.5) $L_{t}=\sum_{k=1}^{\infty}\sqrt{\lambda_{k}}e_{k}l_{t}^{k},\quad t\in[0,T],$
where $e_{k}\in V$, $k\in\mathbb{N}$ are elements of a complete orthonormal system of $H$, $(l^{k}_{t})_{t\in[0,T]}$ are independent and identically distributed (i.i.d.) one-dimensional symmetric Lévy processes with finite second moments, such that $l^{1}_{1}$ has characteristic triplet $(1,\pi_{0},0)$, and $\lambda_{k}\geq 0$ with $\{\lambda_{k}\}\in\ell^{1}$, and where the series converges weakly in $L^{2}(\Omega,{\mathcal{F}},\mathbb{P};H)$ for every $t\in[0,T]$.
Moreover, we assume that there exist $\sigma_{k}\geq c_{0}^{2}>0$, $k\in\mathbb{N}$, with
(3.6) $\|e_{k}\|_{V}\leq\sqrt{\sigma_{k}},\quad k\in\mathbb{N},$
such that
$\{\sigma_{k}\lambda_{k}\}\in\ell^{1}$
@@
[…]
Remark 3.3.
  1. (i)
    Hypothesis (G) is needed to verify a small ball probability for the noise in $V$, see Appendix C, which is only used in the coupling estimate Lemma 5.8 below. When $V$ is a separable Hilbert space, Parseval’s identity yields that for verifying Hypothesis (G) it is sufficient to assume that $\{\sqrt{\sigma_{k}}e_{k}\}$ is a complete orthonormal system of $V$. In this case $\|e_{k}\|_{V}=\sqrt{\sigma_{k}}$ for $k\in\mathbb{N}$. See [6, Theorem 4.7] for cylindrical representations of Lévy processes of the above type.
  2. (ii)
    Note that $\{\lambda_{k}\}\in\ell^{1}$ implies $\{\sqrt{\lambda_{k}}\}\in\ell^{2}$. Furthermore, if $G\equiv 0$, by Hypothesis (G), the operator $B:H\to H$, defined by
    $Bv:=\sum_{k=1}^{\infty}\sqrt{\lambda_{k}}\langle v,e_{k}\rangle e_{k},$
    has finite Hilbert-Schmidt norm
    $\|B\|_{L_{2}(H)}=\left(\sum_{k=1}^{\infty}\|Be_{k}\|^{2}_{H}\right)^{1/2}=\|\{\sqrt{\lambda_{k}}\}\|_{\ell^{2}}<\infty.$
    Hence, in this case, Hypothesis (G) implies Hypothesis (F).
@@ 3.3. Existence and uniqueness of solutions
An $H$-valued càdlàg [^1] That is, right-continuous with left limits. $({\mathcal{F}}_{t})_{t\in[0,T]}$-adapted process $(X_{t})_{t\in[0,T]}$ is called a solution to (2.1) if for its $\mathrm{d}t\otimes\mathbb{P}$-equivalence class $\widehat{X}$, we have
@@
[…]
Now, we recall the existence and uniqueness result given in [14], and [50], respectively. Compare also [3] for a novel alternative approach to locally monotone SPDEs using interpolation spaces.
Proposition 3.5 (Existence and uniqueness).
Let $T>0$ be fixed. Assume that Hypotheses (A), (B), (C), (D) and (F) hold true. Then for any $x\in L^{\beta+2}(\Omega,{\mathcal{F}}_{0},\mathbb{P};H)$, there exists a unique solution to (2.1) in the sense of Definition 3.4.
@@
By our hypotheses, the conditions of [14, Theorem 1.2] are satisfied. ∎
@@ 3.4. Invariant measures
where $B_{b}(H):=\{F:H\to\mathbb{R}:\,F\textrm{ is bounded and Borel measurable}\}$. See [56, Proposition 4.3.5] for a proof of the Markov property in the Gaussian noise case. The Lévy noise case follows by taking essentially the same steps. See also [34, Section 6.4] for a discussion of the Markov property for additive noise SPDEs. For a semigroup $(P_{t})_{t\geq 0}$, we define the dual semigroup $(P_{t}^{\ast})_{t\geq 0}$ acting on ${\mathcal{M}}_{1}(H,{\mathcal{B}}(H)):=\{\mu:{\mathcal{B}}(H)\to[0,1]:\,\mu\textrm{ is a probability measure}\}$ by
$P_{t}^{\ast}\mu(A):=\int_{H}P_{t}1_{A}(x)\,\mu(\mathrm{d}x)\quad\textrm{ for any }\quad A\in{\mathcal{B}}(H),$
where ${\mathcal{B}}(H)$ denote the Borel sets of $H$ and $1_{A}$ denotes the indicator function of the set $A$, see [20] for details.
@@
[…]
@@
$\mathbb{P}_{\mu}\text{-}\lim_{T\to+\infty}\frac{1}{T}\int_{0}^{T}F(X_{t})\,\mathrm{d}t=\int_{H}F\,\mathrm{d}\mu_{\ast},$
where $\mu_{\ast}$ denotes the invariant measure of $\{P_{t}\}_{t\geq 0}$ and $\{X_{t}\}_{t\geq 0}$ denotes the Markov process related to $\{P_{t}\}_{t\geq 0}$ whose initial distribution is $\mu$ and whose path measure is $\mathbb{P}_{\mu}$, and where the convergence takes place in $\mathbb{P}_{\mu}$-probability.
As noted in [46, Remark 3], (3.8) implies uniqueness of the invariant measure. Let us recall [46, Theorem 2], which is an extension of the so-called lower bound technique by Lasota and Szarek [52]. Define
@@
Then there exists a unique invariant measure $\mu_{\ast}$ for $(P_{t})_{t\geq 0}$ such that the semigroup $(P_{t})_{t\geq 0}$ is weak-mean ergodic and the weak law of large numbers holds.
@@ 4. Examples
As $V$ is reflexive, the proof of [34, Proposition 7.1] can be adapted verbatim. ∎
@@
We note that the results of this section also hold in an analog manner if one replaces the bounded domain ${\mathcal{O}}\subset\mathbb{R}^{d}$ with Dirichlet boundary conditions with the flat torus $\mathbb{T}^{d}:=\mathbb{R}^{d}/(2\pi\mathbb{Z}^{d})$, that is, periodic boundary conditions, combined with the requirement that $\int_{\mathbb{T}^{d}}X_{t}\,\mathrm{d}x=0$, $\mathbb{P}$-a.e. for every $t\geq 0$.
@@ 4.1. Stochastic heat equation
Let us prove Theorem 2.2. Consider the stochastic heat equation on a bounded domain ${\mathcal{O}}\subset\mathbb{R}^{d}$ with Lipschitz boundary,
(4.4) $\mathrm{d}X_{t}=\Delta X_{t}\,\mathrm{d}t+B\,\mathrm{d}W_{t}+\int_{D^{c}}G(z)\,\widetilde{N}(\mathrm{d}t,\mathrm{d}z),\quad X_{0}=x,$
where $x\in H:=L^{2}({\mathcal{O}})$. We impose Dirichlet boundary conditions, so that $V:=W^{1,2}_{0}({\mathcal{O}})$. Due to Poincaré’s inequality, we shall equip $V$ with the norm
@@
The hemicontinuity has been proved in [67, Example 4.1.7].
@@
The coercivity has been proved in [67, Example 4.1.7], with $\alpha=2$ and $\delta_{1}=2$.
@@
The monotonicity has been proved in [67, Example 4.1.7] with $\alpha=2$, $\beta=0$, $\delta_{2}=2$, $\rho\equiv 0$, and $C_{2}=0$.
@@
The growth condition has been proved in [67, Example 4.1.7].
@@
Hence, Theorem 2.2 follows immediately from Theorem 2.1.
@@ 4.2. Semilinear stochastic equations
Let us prove Theorem 2.3. Consider the stochastic semilinear equation on a bounded domain ${\mathcal{O}}\subset\mathbb{R}^{d}$, $d\in\mathbb{N}$, $\nu>0$, with Lipschitz boundary,
(4.5) $\mathrm{d}X_{t}=\nu\Delta X_{t}\,\mathrm{d}t+\sum_{i=1}^{d}f_{i}(X_{t})\partial_{i}X_{t}\,\mathrm{d}t+g(X_{t})\,\mathrm{d}t+B\,\mathrm{d}W_{t}+\int_{D^{c}}G(z)\,\widetilde{N}(\mathrm{d}t,\mathrm{d}z),\quad X_{0}=x,$
where $x\in H:=L^{2}({\mathcal{O}})$. We impose Dirichlet boundary conditions, so that $V:=W^{1,2}_{0}({\mathcal{O}})$. Due to Poincaré’s inequality, we shall equip $V$ with the norm
@@
Note that the embedding constant $c_{0}$ is the inverse of the Poincaré constant of ${\mathcal{O}}$. We assume that $f_{i}\in\operatorname{Lip}_{b}(\mathbb{R})$, $i=1,\ldots,d$ for $d=1,2$. For $d=3$, we consider the case that $f_{i}:=b_{i}\in L^{\infty}({\mathcal{O}})$, $i=1,\ldots,3$. Denote $\mathbf{f}:=(f_{1},\ldots,f_{d})^{\textup{t}}$. We assume that $g:\mathbb{R}\to\mathbb{R}$ is continuous with $g(0)=0$, and that there exist $s,C,c\geq 0$, such that
@@
Assume for $d=3$ that $s=\frac{4}{3}$. For $d=1$, assume that $c<\nu c_{0}^{2}$. For $d=2,3$, assume that
@@
The hemicontinuity can be proved as in [56, Lemma 5.1.6 and Example 5.1.7].
@@
The coercivity can be proved as in [56, Lemma 5.1.6 and Example 5.1.7] with $\alpha=2$.
@@
The monotonicity for $\alpha=2$ can be proved as in [56, Lemma 5.1.6 and Example 5.1.7]. As we have to give some attention to the constant $\delta_{2}>0$, we will discuss part of the arguments here. Let $d=1$ and $c<\nu c_{0}^{2}$. Then, by [56, Equation (5.15) and Example 5.1.8],
@@
Let $d=2$. Then, by [56, Equations (5.15) and (5.16), and Example 5.1.8],
@@
Let $d=3$. Then by [56, Lemma 5.1.6 and Example 5.1.8], and taking into account the interpolation inequality
$\|u\|_{L^{4}}^{2}\leq 2\sqrt{2}\|u\|_{H}^{\frac{1}{2}}\|u\|_{V}^{\frac{3}{2}},\quad u\in W_{0}^{1,2}({\mathcal{O}}),$
this case follows by similar arguments, with $\alpha=2$ and $\beta=\frac{10}{3}$, see [56, Example 5.1.8 (3)].
@@
Hypothesis (D) follows as in [56, Lemma 5.1.6 and Example 5.1.7].
@@
Consequently, Theorem 2.3 follows directly from Theorem 2.1.

### v4

@@ 3.2. Hypotheses on the noise
The noise coefficient $B:U\to H$ is linear, bounded and satisfies
(3.3) $B\in L_{2}(U,H).$
Hypothesis (G) ($V$-Regularity of the noise).
For any $T>0$, the driving process $\{B\,W_{t}\}_{t\in[0,T]}$ satisfies a small ball property in $V$, that is, for any $\delta>0$,
(3.4) $\mathbb{P}\left(\sup_{t\in[0,T]}\|B\,W_{t}\|_{V}\leq\delta\right)>0.$
Remark 3.3.
If $V$ is a separable Hilbert space, Hypothesis (G) follows from
(3.5) $B\in L_{2}(U,V).$
This can be easily verified by standard properties of Brownian motion by choosing a cylindrical representation.
3.3. Hypotheses for ergodicity and mixing
Hypothesis (H) (Ergodicity).
Assume that $0\leq\beta\leq\alpha-2$. Assume that $\eta\equiv 0$ or $\rho\equiv 0$ (locally monotone case). Furthermore, for $\lambda_{0}\in(0,1)$ and $\lambda_{i}\in(0,1)$, $i=1,2,3$, such that $\sum_{i=0}^{3}\lambda_{i}=1$, and for
(3.6) $\begin{split}&c_{1}:=\frac{\alpha(\beta+2)}{2(\alpha+\beta)\left(\lambda_{1}\delta_{1}c_{0}^{\alpha}\frac{\alpha+\beta}{\beta}\right)^{\frac{\beta}{\alpha}}}\|B\|_{L_{2}(U,H)}^{\frac{2(\alpha+\beta)}{\alpha}},\\ &c_{2}:=\beta^{\frac{\alpha+\beta}{\alpha}}\frac{\alpha(\beta+2)}{(\alpha+\beta)\left(\lambda_{2}\delta_{1}c_{0}^{\alpha}\frac{\alpha+\beta}{2\beta}\right)^{\frac{\beta}{\alpha}}}\|B\|_{L(U,H)}^{\frac{2(\alpha+\beta)}{\alpha}},,\\ &c_{3}:=\frac{(\beta+2)(\alpha-\beta-2)}{2(\alpha+\beta)\left(\lambda_{3}\delta_{1}c_{0}^{\alpha}\frac{\alpha+\beta}{2\beta+2}\right)^{\frac{2\beta+2}{\alpha-\beta-2}}}(\beta+2)^{\frac{\alpha+\beta}{\alpha-\beta-2}}\|B\|_{L_{2}(U,H)}^{\frac{2(\alpha+\beta)}{\alpha-\beta-2}},\end{split}$
we assume that
(3.7) $\frac{2(c_{1}+c_{2}+c_{3})C_{2}}{\lambda_{0}\delta_{1}(\beta+2)}\leq\delta_{2},$
where $\delta_{2}>0$ and $C_{2}\geq 0$ are as in Hypothesis (C). Furthermore, if $\alpha=\beta+2$, then $c_{3}=0$ and we assume additionally that
(3.8) $\|B\|^{2}_{L_{2}(U,H)}\leq\lambda_{3}\frac{1}{\alpha}\delta_{1}c_{0}^{\alpha}.$
Hypothesis (I) (Exponential ergodicity and mixing).
Assume that $0\leq\beta\leq\alpha-2$. Assume that $\eta\equiv 0$ or $\rho\equiv 0$ (locally monotone case). In the situation of Hypothesis (H), assume that there exists $\gamma\in(0,\delta_{2}]$, such that
(3.9) $\frac{2(c_{1}+c_{2}+c_{3})C_{2}}{\lambda_{0}\delta_{1}(\beta+2)}\leq\delta_{2}-\gamma.$
Remark 3.4.
If $\beta=0$, then $\lambda_{1}=\lambda_{2}=0$ and
$c_{1}=\|B\|^{2}_{L_{2}(U,H)},$
$c_{2}=0.$
For $\alpha=2$ and $\beta=0$,
$c_{1}=\|B\|^{2}_{L_{2}(U,H)},$
$c_{2}=0$
$c_{3}=0,$
and Hypothesis (H) simplifies to $\lambda_{0}=1-\lambda_{3}$,
(3.10) $\frac{2\|B\|^{2}_{L_{2}(U,H)}C_{2}}{\lambda_{0}\delta_{1}}\leq\delta_{2},$
@@ 3.3. Hypotheses for ergodicity and mixing
(3.11) $\|B\|^{2}_{L_{2}(U,H)}\leq\frac{1}{2}\lambda_{3}\delta_{1}c_{0}^{2}.$
Accordingly, Hypothesis (I) simplifies to $\gamma\in(0,\delta_{2}]$,
(3.12) $\frac{2\|B\|^{2}_{L_{2}(U,H)}C_{2}}{\lambda_{0}\delta_{1}}\leq\delta_{2}-\gamma,$
@@
(3.13) $\|B\|^{2}_{L_{2}(U,H)}\leq\frac{1}{2}\lambda_{3}\delta_{1}c_{0}^{2}.$
@@ 3.4. Existence and uniqueness of solutions
An $H$-valued continuous $({\mathcal{F}}_{t})_{t\in[0,T]}$-adapted process $(X_{t})_{t\in[0,T]}$ is called a solution to (2.1) if for its $\mathrm{d}t\otimes\mathbb{P}$-equivalence class $\widehat{X}$, we have
@@
    $X_{t}=x+\int_{0}^{t}A(\overline{X}_{s})\,\mathrm{d}s+B\,W_{t},\quad t\in[0,T],\quad\mathbb{P}\text{-a.s.}$
Now, we recall the existence and uniqueness result given in [70]. Compare with [16] for Lévy noise. Compare also [4] for a novel alternative approach to locally monotone SPDEs using interpolation spaces.
Proposition 3.6 (Existence and uniqueness (locally monotone case)).
Let $T>0$ be fixed. Assume that Hypotheses (A), (B), (C), (D) and (F) hold true. Assume that $\eta\equiv 0$ or $\rho\equiv 0$ (locally monotone case). Then for any $x\in L^{\beta+2}(\Omega,{\mathcal{F}}_{0},\mathbb{P};H)$, there exists a unique solution to (2.1) in the sense of Definition 3.5.
@@
By our hypotheses, the conditions of [70, Theorem 1.2] are satisfied. ∎
For the fully locally monotone case, we obtain probabilistically weak solutions, see [83, 63].
Proposition 3.7 (Existence and uniqueness (fully locally monotone case)).
Let $T>0$ be fixed. Assume that Hypotheses (A), (B), (C), (D) and (F) hold true. Assume that both $\eta$ and $\rho$ are non-trivial (fully locally monotone case). Assume that $V\subset H$ compactly. Then there exists a stochastic basis $(\tilde{\Omega},\tilde{{\mathcal{F}}},\{\tilde{{\mathcal{F}}}_{t}\}_{t\geq 0},\tilde{\mathbb{P}},\{\tilde{W}_{t}\}_{t\geq 0})$ such that for any $x\in H$, there exists a unique probabilistically weak solution to (2.1) in the sense of the statements of Definition 3.5 holding with respect to this stochastic basis.
Proof.
By our hypotheses, the conditions of [83, Theorem 2.6] are satisfied. ∎
@@ 3.5. Invariant measures
where $B_{b}(H):=\{F:H\to\mathbb{R}:\,F\textrm{ is bounded and Borel measurable}\}$. See [70, Proposition 4.3.5] for a proof of the Markov property in the Gaussian noise case. See also [42, Section 6.4] for a discussion of the Markov property for additive noise SPDEs. For a semigroup $(P_{t})_{t\geq 0}$, we define the dual semigroup $(P_{t}^{\ast})_{t\geq 0}$ acting on ${\mathcal{M}}_{1}(H,{\mathcal{B}}(H)):=\{\mu:{\mathcal{B}}(H)\to[0,1]:\,\mu\textrm{ is a probability measure}\}$ by
$P_{t}^{\ast}\mu(A):=\int_{H}P_{t}\mathbbm{1}_{A}(x)\,\mu(\mathrm{d}x)\quad\textrm{ for any }\quad A\in{\mathcal{B}}(H),$
where ${\mathcal{B}}(H)$ denote the Borel sets of $H$ and $\mathbbm{1}_{A}$ denotes the indicator function of the set $A$, see [24] for details.
@@
Note that, under certain conditions, the $e$-property can be derived from the eventual $e$-property [72].
@@
(3.14) $\text{w-}\lim_{T\to\infty}\frac{1}{T}\int_{0}^{T}P_{t}^{\ast}\mu\,\mathrm{d}t=\mu_{\ast}\quad\text{for every }\quad\mu\in{\mathcal{M}}_{1}(H,{\mathcal{B}}(H)),$
@@
$\mathbb{P}_{\mu}\text{-}\lim_{T\to\infty}\frac{1}{T}\int_{0}^{T}F(X_{t}^{\mu})\,\mathrm{d}t=\int_{H}F\,\mathrm{d}\mu_{\ast},$
where $\mu_{\ast}$ denotes the invariant probability measure of $\{P_{t}\}_{t\geq 0}$ and $\{X_{t}^{\mu}\}_{t\geq 0}$ denotes the Markov process related to $\{P_{t}\}_{t\geq 0}$ whose initial distribution is $\mu$ and whose path measure is $\mathbb{P}_{\mu}$, and where the convergence takes place in $\mathbb{P}_{\mu}$-probability.
As noted in [57, Remark 3], (3.14) implies uniqueness of the invariant probability measure. Let us recall [57, Theorem 2], which is an extension of the lower bound technique by Lasota and Szarek [65]. Define
@@
Then there exists a unique invariant probability measure $\mu_{\ast}$ for $(P_{t})_{t\geq 0}$ such that the semigroup $(P_{t})_{t\geq 0}$ is weak-mean ergodic and the weak law of large numbers holds.
@@ 4. Examples
As $V$ is reflexive, the proof of [42, Proposition 7.1] can be adapted verbatim. ∎
@@
We note that the results of this section also hold in an analog manner if one replaces the domain of finite width ${\mathcal{O}}\subset\mathbb{R}^{d}$ with Dirichlet boundary conditions with the flat torus $\mathbb{T}^{d}:=\mathbb{R}^{d}/(2\pi\mathbb{Z}^{d})$, that is, periodic boundary conditions, combined with the requirement that $\int_{\mathbb{T}^{d}}X_{t}\,\mathrm{d}x=0$, $\mathbb{P}$-a.e. for every $t\geq 0$.
@@ 4.1. Stochastic heat equation
Let us prove Theorem 2.3. Consider the stochastic heat equation for $\nu>0$ on a domain of finite width ${\mathcal{O}}\subset\mathbb{R}^{d}$ with Lipschitz boundary,
(4.4) $\mathrm{d}X_{t}=\nu\Delta X_{t}\,\mathrm{d}t+B\,\mathrm{d}W_{t},\quad X_{0}=x,$
where $x\in H:=L^{2}({\mathcal{O}})$. We impose Dirichlet boundary conditions, so that $V:=W^{1,2}_{0}({\mathcal{O}})$. Due to Poincaré’s inequality [1], we shall equip $V$ with the norm
@@
The hemicontinuity has been proved in [82, Example 4.1.7].
@@
The coercivity has been proved in [82, Example 4.1.7], with $\alpha=2$ and $\delta_{1}=2\nu$.
@@
The monotonicity has been proved in [82, Example 4.1.7] with $\alpha=2$, $\beta=0$, $\delta_{2}=2\nu c_{0}^{2}$, $\rho\equiv 0$, and $C_{2}=0$.
@@
The growth condition has been proved in [82, Example 4.1.7].
@@
Hence, Theorem 2.3 follows immediately from Theorem 2.1 and Theorem 6.2.
@@ 4.2. Semilinear stochastic equations
Let us prove Theorem 2.4. Consider the stochastic semilinear equation on a domain of finite width ${\mathcal{O}}\subset\mathbb{R}^{d}$, $d\in\mathbb{N}$, $\nu>0$, with Lipschitz boundary,
(4.5) $\mathrm{d}X_{t}=\nu\Delta X_{t}\,\mathrm{d}t+\sum_{i=1}^{d}f_{i}(X_{t})\partial_{i}X_{t}\,\mathrm{d}t+g(X_{t})\,\mathrm{d}t+B\,\mathrm{d}W_{t}\quad X_{0}=x,$
where $x\in H:=L^{2}({\mathcal{O}})$. We impose Dirichlet boundary conditions, so that $V:=W^{1,2}_{0}({\mathcal{O}})$. Due to Poincaré’s inequality [1], we shall equip $V$ with the norm
@@
Note that the embedding constant $c_{0}$ is the inverse of the Poincaré constant of $\mathcal{O}$}. We assume that $f(x)=x$ in $d=1$ or that $f_{i}\in\operatorname{Lip}_{b}(\mathbb{R})$, $i=1,\ldots,d$ for $d=1,2$. Denote $\mathbf{f}:=(f_{1},\ldots,f_{d})^{\textup{t}}$. We assume that $g:\mathbb{R}\to\mathbb{R}$ is continuous with $g(0)=0$, and that there exist $C,c\geq 0$, $s\in[0,2]$, such that
@@
For $d=1$, assume that $c<\nu c_{0}^{2}$. For $d=2$, assume that
@@
The hemicontinuity can be proved as in [70, Lemma 5.1.6 and Example 5.1.7].
@@
The coercivity can be proved as in [70, Lemma 5.1.6 and Example 5.1.7] with $\alpha=2$.
@@
The monotonicity for $\alpha=2$ can be proved as in [70, Lemma 5.1.6 and Example 5.1.7]. As we have to give some attention to the constant $\delta_{2}>0$, we will discuss part of the arguments here. Let $d=1$ and $c<\nu c_{0}^{2}$. Then, by [70, Equation (5.15) and Example 5.1.8],
@@
Let $d=2$. Then, by [70, Equations (5.15) and (5.16), and Example 5.1.8],
@@
Hypothesis (D) follows as in [70, Lemma 5.1.6 and Example 5.1.7].
@@
If $d=1$ and $f(x)=x$, then by integration by parts,
$\langle(\mathbf{f}(u)\cdot\nabla u),u\rangle=\langle u\partial_{x}u,u\rangle=-2\langle u\partial_{x}u,u\rangle=0.$
@@
Consequently, Theorem 2.4 follows directly from Theorem 2.1 and Theorem 6.2.